find the area bounded by y=|e^x-1| of x axis in the interval [-1,1]

Grade:12

1 Answers

bharat bajaj IIT Delhi
10 years ago
y=|e^x-1|
Check when e^x - 1 is positive and negative,
So, y= -(e^x - 1) -1<x<0
y= +(e^x - 1) 0<x<1
Now integrate,
Integral(-1 to 0) of -
(e^x - 1) = +1/e
Integral(0 to 1) of
(e^x - 1) = e
Area bounded = e + 1/e

Thanks & Regards
Bharat Bajaj
askIITians Faculty
IIT Delhi